OpenStudy (gabylovesyou):

@aum The volume of a cube is 36 cubic feet. What is the volume of a pyramid that fits exactly inside the cube? Provide an explanation and proof for your answer to receive full credit.

OpenStudy (campbell_st):

Volume of a Pyramid is V = 1/3 length x width x height Volume of a cube is V = length x width x height hopefully you can see the connection

OpenStudy (gabylovesyou):

so its 12 cubic feet ? @campbell_st

OpenStudy (anonymous):

volume of pyramid = 1/3 of cube =(1/3)36 =12 cubic feet

OpenStudy (aum):

Yes!

OpenStudy (gabylovesyou):

Thanks everyone! :D

OpenStudy (campbell_st):

that's correct... well done

OpenStudy (gabylovesyou):

why 1/3 tho ? dont 6 pyramids fit in a cube ?

OpenStudy (campbell_st):

basically 3 pyramids fit in a cube with the same dimensions... in a similar way 3 cones fit in a cylinder with the same radius and height

OpenStudy (gabylovesyou):

oh ok thanks